# TransNAS-TSAD: Harnessing Transformers for Multi-Objective Neural Architecture Search in Time Series Anomaly Detection [^1] This document is the results of the research project funded by the National Science Foundation.22footnotemark: 2

### 03 Dec 2023 v2

@@ 4.1 Datasets
Dataset Train Size Test Size Dimensions Sequences Anomalies (%)
NAB 4,033 4,033 1 6 0.92
UCR 1,600 5,900 1 4 1.88
MBA 100,000 100,000 2 8 0.14
SMAP 135,183 427,617 25 55 13.13
MSL 58,317 73,729 55 3 10.72
SWaT 496,800 449,919 51 1 11.98
WADI 1,048,571 172,801 123 1 5.99
SMD 708,405 708,420 38 4 4.16
@@
Our study employs a comprehensive suite of datasets, each meticulously chosen to challenge and validate the robustness of time series anomaly detection methods across a variety of real-world scenarios. Table 2 summarizes the statistics of the datasets. These datasets, renowned for their complexity and diversity, allow for an effective benchmarking of our TransNAS-TSAD approach against existing methodologies.
  • 1.
@@
    Numenta Anomaly Benchmark (NAB) Dataset: OThis dataset encompasses a wide range of real-world data, including temperature sensor readings, cloud machine CPU utilization, service request latencies, and taxi demand data in New York City. It’s important to note some anomalies in labeling within this dataset, particularly in the nyc-taxi trace, which is often excluded from analyses (Ahmad et al., (2017); Nakamura et al., (2020)).
  • 2.
@@
    HexagonML (UCR) Dataset: Featured in the KDD 2021 cup, this dataset comprises a diverse collection of univariate time series. In our research, we selectively utilized datasets that are derived from real-world sources, focusing specifically on the InternalBleeding and ECG datasets. This choice was driven by our need for natural data representations in our analysis. Consequently, we did not include any synthetic sequences that are also part of the HexagonML (UCR) Dataset (Dau et al., (2018)).”
  • 3.
@@
    MIT-BIH Supraventricular Arrhythmia Database (MBA) Dataset: This dataset includes electrocardiogram recordings from four patients, featuring anomalies such as supraventricular contractions or premature heartbeats. It’s a well-recognized dataset in data management studies (Moody & Mark, (2001); Boniol et al., (2020)).
  • 4.
    Soil Moisture Active Passive (SMAP) Dataset:: Provided by NASA, this dataset consists of soil samples and telemetry data from the Mars rover (Hundman et al., (2018)).
  • 5.
@@
    Mars Science Laboratory (MSL) Dataset: Similar to the SMAP dataset, the MSL dataset includes sensor and actuator data from the Mars rover.(Hundman et al., (2018)) Due to the presence of many trivial sequences, only specific non-trivial sequences are typically analyzed such as (A4, C2 and T1) pointed by (Nakamura et al., (2020); Tuli et al., (2022)).
  • 6.
@@
    Secure Water Treatment (SWaT) Dataset: This dataset is derived from a real-world water treatment plant, including data from 7 days of normal operations and 4 days under abnormal conditions. It features readings from various sensors and actuators (Mathur & Tippenhauer, (2016)).
  • 7.
@@
    Water Distribution (WADI) Dataset: An expansion of the SWaT dataset, WADI includes a larger array of sensors and actuators. The dataset spans over a longer period, with 14 days of normal operation and 2 days under attack scenarios (Ahmed et al., (2017)).
  • 8.
@@
    Server Machine Dataset (SMD): Covering five weeks of data, this dataset includes resource utilization traces from 28 machines in a compute cluster. Only specific non-trivial sequences are used for analysis (Su et al., (2019)).
Each dataset is selected for its potential to challenge and validate the robustness of time series anomaly detection methods across various industry scenarios.
@@ 4.2 Evaluation Criteria
In our study, aimed at advancing anomaly detection in time series data, we employ key metrics aligned with the objectives of the TransNAS-TSAD framework. These metrics are selected to optimize anomaly detection effectiveness, considering practical deployment aspects.
4.2.1 F1 Score as Primary Performance Indicator
The F1 score, the harmonic mean of precision and recall, is central to our evaluation strategy. It is particularly effective in addressing the challenge of imbalanced datasets common in anomaly detection, offering a balanced performance measure by accounting for the trade-off between false positives and false negatives.
@@ 4.2.2 Introducing the Efficiency-Accuracy-Complexity Score (EACS)
Model complexity is often indicated by parameter count. However, this metric can be misleading, especially for transformer-based models like TransNAS-TSAD. These models, despite their high parameter count, capture complex data patterns effectively, justifying their complexity. Therefore, we report parameter count and interpret it in the context of the model’s advanced capabilities. To comprehensively evaluate TransNAS-TSAD, we introduce the Efficiency-Accuracy-Complexity Score (EACS), a composite metric encompassing accuracy, training efficiency, and model complexity. It is defined as:
$EACS=w_{\text{acc}}\times\left(\frac{\text{F1 Score}}{\text{Max F1 Score}}\right)+w_{\text{time}}\times\left(\frac{\text{Training Time}}{\text{Max Training Time}}\right)+w_{\text{params}}\times\left(\frac{\text{Parameter Count}}{\text{Max Parameter Count}}\right)$
where $w_{acc}$, $w_{time}$, and $w_{params}$ are weights for the normalized F1 score, inverse training time, and inverse parameter count, respectively. Here, the weights are set as $w_{acc}=0.4$, $w_{time}=0.4$, and $w_{params}=0.2$, balancing accuracy, training efficiency, and model simplicity. This weighting scheme aligns with our research goals and practical deployment considerations, prioritizing accuracy and training efficiency while considering model complexity.
Transformer models, known for their high parameter count, offer efficient processing capabilities through parallelization, making them suitable for scenarios requiring high accuracy where computational resources are not a constraint.
4.3 Note on Additional Metrics
Other metrics, such as ROC/AUC, are also valuable; however, our focus on the F1 score, parameter count, and EACS aligns with the specific goals of TransNAS-TSAD. This approach ensures a balanced evaluation of the models, considering both their effectiveness in anomaly detection and practical deployability.
@@ 5 Results
5.1 Evaluating TransNAS-TSAD: Precision, Recall, and F1 Score Analysis
In our detailed evaluation of anomaly detection models, as captured in Table 3, the TransNAS-TSAD model emerges as a formidable contender, showcasing its prowess across diverse datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD, with metrics of precision, recall, and F1 scores serving as the benchmarks.
Standing out for its exceptional performance, TransNAS-TSAD notably surpasses other models in datasets such as NAB, UCR, MBA, and SMAP, with remarkable F1 scores (0.9411, 0.9910, 0.9861, and 0.9510 respectively), thereby demonstrating its adeptness in handling both univariate and multivariate time series. This is particularly evident when compared to models like OmniAnomaly and MSCRED, where TransNAS-TSAD’s ability to navigate complex data patterns becomes apparent.
However, in the MSL dataset, the TranAD model, with its nuanced approach, shows a marginally superior performance with an F1 score of 0.9494 compared to TransNAS-TSAD’s 0.9482, possibly indicative of TranAD’s specific strengths in handling the dataset’s characteristics. Despite this, TransNAS-TSAD regains its lead in the SWaT dataset with an F1 score of 0.8314, highlighting its adaptability to various industrial contexts.
Its exemplary performance continues in the challenging WADI and SMD datasets, where it achieves the highest F1 scores of 0.7735 and 0.9986, respectively, showcasing its unrivaled efficiency in detecting subtle anomalies in complex, high-dimensional datasets. This overarching success of TransNAS-TSAD, especially when contrasted with TranAD’s performance, underscores its comprehensive capability in anomaly detection.
While models like LSTM-NDT and DAGMM show variable performance across different datasets, their inconsistency underlines TransNAS-TSAD’s steady effectiveness across various scenarios. Similarly, while OmniAnomaly and MSCRED perform well in certain contexts, they fall short of TransNAS-TSAD’s sophisticated temporal pattern handling.
Other competitive models, including MAD-GAN and USAD, are eclipsed by TransNAS-TSAD’s more holistic approach, and even though CAE-M and GDN demonstrate strengths in specific scenarios, they generally do not match TransNAS-TSAD’s overall adaptability.
The superiority of TransNAS-TSAD, especially in comparison to TranAD, invites further investigation through critical difference analysis, which would reinforce the statistical significance of the observed performance differences and provide a deeper understanding of TransNAS-TSAD’s robustness and reliability against a backdrop of diverse and challenging anomaly detection scenarios.
Method NAB UCR MBA SMAP
P R F1 P R F1 P R F1 P R F1
LSTM-NDT 0.6400 0.6667 0.6531 0.5231 0.8294 0.5231 0.9207 0.9718 0.9456 0.8523 0.7326 0.7879
DAGMM 0.7622 0.7292 0.7453 0.5337 0.9718 0.5337 0.9475 0.9900 0.9683 0.8069 0.9891 0.8888
OmniAnomaly 0.8421 0.6667 0.7442 0.8346 0.9999 0.8346 0.8561 1.000 0.9225 0.813 0.9419 0.8728
MSCRED 0.8522 0.6700 0.7502 0.5441 0.9718 0.5441 0.9272 1.0000 0.9623 0.8175 0.9216 0.8664
MAD-GAN 0.8666 0.7012 0.7752 0.8538 0.9891 0.8538 0.9396 1.0000 0.9689 0.8157 0.9216 0.8654
USAD 0.8421 0.6667 0.7442 0.8952 1.0000 0.8952 0.8953 0.9989 0.9443 0.748 0.9627 0.8419
CAE-M 0.7918 0.8019 0.7968 0.6981 1.0000 0.6981 0.8442 0.9997 0.9154 0.8193 0.9567 0.8827
GDN 0.8129 0.7872 0.7998 0.6894 0.9988 0.6894 0.8832 0.9892 0.9332 0.748 0.9891 0.8518
TranAD 0.8889 0.9892 0.9364 0.9407 1.0000 0.9407 0.9569 1.0000 0.978 0.8043 0.9999 0.8915
TransNAS-TSAD 0.8888 1.0000 0.9411 0.9823 1.0000 0.9910 0.9726 1.0000 0.9861 0.9066 1.0000 0.9510
Method MSL SWaT WADI SMD
P R F1 P R F1 P R F1 P R F1
LSTM-NDT 0.6288 1.0000 0.7721 0.7778 0.5109 0.6167 0.0138 0.7823 0.0271 0.9736 0.844 0.9042
DAGMM 0.7363 1.0000 0.8482 0.9933 0.6879 0.8128 0.076 0.9981 0.1412 0.9103 0.9914 0.9491
OmniAnomaly 0.7848 0.9924 0.8765 0.9782 0.6957 0.8131 0.3158 0.6541 0.426 0.8881 0.9985 0.9401
MSCRED 0.8912 0.9862 0.9363 0.9992 0.677 0.8072 0.2513 0.7319 0.3741 0.7276 0.9974 0.8414
MAD-GAN 0.8516 0.993 0.9169 0.9593 0.6957 0.8065 0.2233 0.9124 0.3588 0.9991 0.844 0.915
USAD 0.7949 0.9912 0.8822 0.9977 0.6879 0.8143 0.1873 0.8296 0.3056 0.906 0.9974 0.9495
CAE-M 0.7751 1.0000 0.8733 0.9697 0.6957 0.8101 0.2782 0.7918 0.4117 0.9082 0.9671 0.9367
GDN 0.9308 0.9892 0.9591 0.9697 0.6957 0.8101 0.2912 0.7931 0.426 0.717 0.9974 0.8342
TranAD 0.9038 0.9999 0.9494 0.976 0.6997 0.8151 0.3529 0.8296 0.4951 0.9262 0.9974 0.9605
TransNAS-TSAD 0.9567 1.000 0.94.82 0.9415 0.7624 0.8314 0.8626 0.7340 0.7735 0.9985 0.9988 0.9986
Table 3: Benchmarking TransNAS-TSAD against Baseline Models: A summary of precision (P), recall (R), and F1 scores showcases the advanced anomaly detection capabilities of TransNAS-TSAD, derived via NSGA-II-based NAS, across diverse datasets.
@@ 5.1 Evaluating TransNAS-TSAD: Precision, Recall, and F1 Score Analysis
5.2 In-Depth Analysis of Training Efficiency and EACS in Anomaly Detection Models
Furthering our investigation from Table 3, Table 4 provides a more granular look at the training efficiency and overall performance of anomaly detection models, with a particular emphasis on the Efficiency-Accuracy-Complexity Score (EACS). This table not only reiterates the superior performance of TransNAS-TSAD in precision and recall metrics, as observed in Table 3 but also introduces a critical analysis of training times and parameter counts across a range of datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD.
The training time, a pivotal yet often overlooked aspect, emerges as a decisive factor in evaluating the practicality of models. In operational settings, where rapid model deployment and iterative updating are essential, training time gains precedence over traditional metrics like parameter count. This is evidenced in the case of TransNAS-TSAD, which, across diverse datasets, demonstrates not only high F1 scores but also markedly reduced training times, leading to its high EACS. In the NAB dataset, for instance, TransNAS-TSAD achieves an F1 score of 94.11% with a training time of just 2.70 sec, resulting in an EACS of 0.9742. This trend of balancing accuracy with training efficiency is replicated in the UCR dataset, where it achieves an F1 score of 99.10% and an even lower training time of 2.24 sec, leading to an EACS of 0.9922.
Diving deeper, in datasets like MBA and SMAP, TransNAS-TSAD’s EACS scores of 0.9932 and 0.9734, respectively, are not solely the product of its accuracy but are significantly bolstered by its training time efficiency. For example, in MBA, its training time of 2.27 is among the lowest, demonstrating the model’s capability to quickly adapt and learn from complex multivariate time series data. In comparison, models such as MSCRED, while exhibiting commendable F1 scores, are penalized in their EACS due to their disproportionately high training times (e.g., 774.99 sec for MBA), highlighting a crucial trade-off between model complexity and operational efficiency.
In more challenging datasets like SWaT and WADI, the importance of training time becomes even more pronounced. TransNAS-TSAD’s ability to maintain high F1 scores while drastically cutting down on training times positions it as a highly effective tool for real-world applications. Its EACS of 0.9297 for SWaT and 0.6977 for WADI, while highlighting its proficiency in these industrial datasets, also reflect the balanced approach of the model, juxtaposing accuracy with swift adaptability.
Furthermore, the analysis of models like GDN and TranAD provides additional context. While GDN shows strong performance in specific datasets like MSL with the highest F1 score (95.91%) and a respectable EACS of 0.9191, it falls behind TransNAS-TSAD in terms of training efficiency. Similarly, TranAD, despite its high F1 scores and relatively lower parameter counts, does not consistently match the training efficiency of TransNAS-TSAD, as seen in datasets like SMD and WADI.
In conclusion, Table 4’s comprehensive analysis underscores the critical role of training time in the practical deployment of anomaly detection models. It establishes that models with lower training times, like TransNAS-TSAD, offer a significant advantage in real-world scenarios, effectively balancing accuracy, efficiency, and complexity. This detailed examination not only affirms the superiority of TransNAS-TSAD in standard performance metrics but also in its operational readiness, making it an exemplary model for efficient and effective anomaly detection in diverse and demanding environments.
Method NAB UCR MBA SMAP
F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
LSTM-NDT 65.31 23.40 1710 0.8424 52.31 11.14 1765 0.7985 94.56 47.80 18641 0.9520 78.79 36.43 29146 0.8459
DAGMM 74.53 64.50 1266 0.8467 53.37 27.50 1214 0.7919 96.83 92.36 2448 0.9394 88.80 26.60 7266 0.9050
OmniAnomaly 74.42 88.70 13717 0.8250 83.46 38.21 13025 0.8863 92.25 197.66 25474 0.8649 87.28 66.77 16813 0.8230
MSCRED 75.02 503.60 1237377 0.3001 54.41 559.60 128655 0.2176 96.23 774.99 2441778 0.3849 86.64 25.70 8237452 0.6982
MAD-GAN 77.52 53.90 838 0.8671 85.38 31.50 822 0.9177 96.89 229.78 1877 0.8688 86.54 41.26 6718 0.8683
USAD 74.42 65.40 1359 0.8455 89.52 54.23 1359 0.9172 94.43 191.26 1609 0.8789 84.19 49.22 7395 0.8439
CAE-M 79.68 33.60 7229 0.8909 69.81 22.10 7365 0.8520 91.54 89.66 15411 0.9186 88.27 212.51 7229 0.5529
GDN 79.98 131.60 574 0.8153 68.94 64.70 566 0.8286 93.32 203.45 1106 0.8682 85.18 104.50 2974 0.7440
TranAD 93.64 3.37 615 0.9718 94.07 1.14 619 0.9745 97.80 5.11 1298 0.9885 89.15 5.62 62271 0.9445
TransNAS-TSAD 94.11 2.70 285 0.9742 99.10 2.24 1690 0.9922 98.61 2.27 593 0.9932 94.60 3.19 39045 0.9734
Method MSL SWaT WADI SMD
F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS F1 Training Time # Params EACS
LSTM-NDT 77.21 37.80 61856 0.8887 66.70 41.50 72854 0.7957 2.71 422.67 133712 0.5737 90.42 460.70 42853 0.9100
DAGMM 84.82 25.66 18756 0.9258 81.28 29.66 16558 0.8748 14.12 246.32 39387 0.6392 94.91 337.90 10516 0.9420
OmniAnomaly 87.65 29.69 39753 0.9348 81.31 40.20 36541 0.8568 42.60 361.60 39753 0.7471 94.01 311.20 22941 0.9412
MSCRED 93.63 40.70 18476728 0.7535 80.72 236.80 14227264 0.3229 37.41 1884.25 75530 0.6435 84.14 349.77 14237356 0.6978
MAD-GAN 91.69 32.64 17446 0.9497 80.65 34.70 15958 0.8638 35.88 321.20 35764 0.7228 91.50 424.60 9903 0.9188
USAD 88.22 29.60 14859 0.9374 81.43 35.81 12502 0.8651 30.56 389.72 32859 0.6983 94.95 328.77 10609 0.9432
CAE-M 87.33 774.60 204687 0.5471 81.01 71.22 194525 0.8010 41.17 7724.92 388905 0.3203 93.67 3606.90 114219 0.5731
GDN 95.91 121.80 153541 0.9191 81.01 83.55 115442 0.7813 42.60 6047.12 287082 0.4245 83.42 1000.40 7255 0.8226
TranAD 94.94 6.88 272181 0.9733 81.53 2.10 31336 0.9221 49.51 177.60 1378173 0.6316 96.05 56.70 135110 0.9760
TransNAS-TSAD 94.82 5.44 236491 0.9739 83.14 1.60 13464 0.9297 77.35 225.60 1752919 0.6977 99.81 52.40 132050 0.9916

### 04 Mar 2024 v3

@@ 4.1 Datasets
Dataset Train Size Test Size Dimensions Sequences Anomalies (%) NAB 4,033 4,033 1 6 0.92 UCR 1,600 5,900 1 4 1.88 MBA 100,000 100,000 2 8 0.14 SMAP 135,183 427,617 25 55 13.13 MSL 58,317 73,729 55 3 10.72 SWaT 496,800 449,919 51 1 11.98 WADI 1,048,571 172,801 123 1 5.99 SMD 708,405 708,420 38 4 4.16
@@
Our study employs suite of datasets selected to challenge and validate the robustness of the time series anomaly detection methods across a variety of real-world scenarios. These datasets, renowned for their complexity and diversity, allow for an effective benchmarking of our TransNAS-TSAD approach against existing methodologies. See Table 2 for summary statistics.
@@
  • Numenta Anomaly Benchmark (NAB) Dataset: This dataset encompasses a wide range of real-world data, including temperature sensor readings, cloud machine CPU utilization, service request latencies, and taxi demand data in New York City. It’s important to note some anomalies in labeling within this dataset are excluded from analyses, particularly in the NYC-taxi traces [1, 54].
@@
  • HexagonML (UCR) Dataset: Featured in the KDD 2021 cup, this dataset comprises a diverse collection of univariate time series. In our research, we selectively use natural data representations derived from real-world sources, focusing specifically on the InternalBleeding and ECG datasets. Consequently, we did not include any synthetic sequences that are also part of the HexagonML (UCR) dataset [20].
@@
  • MIT-BIH Supraventricular Arrhythmia Database (MBA) Dataset: This dataset includes electrocardiogram recordings from four patients, featuring anomalies such as supraventricular contractions or premature heartbeats. It’s a well-recognized dataset in data management studies [53, 9].
  • Soil Moisture Active Passive (SMAP) Dataset: This NASA-provided dataset includes global measurements of soil moisture in the top 5 cm of Earth’s soil surface, collected approximately every three days by the SMAP satellite. It is designed to enhance our understanding of water, carbon, and energy cycles[24].
@@
  • Mars Science Laboratory (MSL) Dataset: Similar to the SMAP dataset, the MSL dataset includes sensor and actuator data from the Mars rover.[24] Due to the presence of many trivial sequences, only specific non-trivial sequences are typically analyzed such as (A4, C2 and T1) pointed by [54, 59].
@@
  • Secure Water Treatment (SWaT) Dataset: This dataset is derived from a real-world water treatment plant, including data from 7 days of normal operations and 4 days under abnormal conditions. It features readings from various sensors and actuators [51].
@@
  • Water Distribution (WADI) Dataset: An expansion of the SWaT dataset, WADI includes a larger array of sensors and actuators. The dataset spans a longer period, with 14 days of normal operation and 2 days under attack scenarios [2].
@@
  • Server Machine Dataset (SMD): Covering five weeks of data, this dataset includes resource utilization traces from 28 machines in a compute cluster. Only specific non-trivial sequences are used for analysis [61].
@@ 4.2 Evaluation Criteria
We employ key metrics aligned with the objectives of the TransNAS-TSAD framework. These metrics are selected to optimize anomaly detection effectiveness, considering practical deployment aspects. The F1 score, defined as the harmonic mean of precision and recall, is central to our evaluation strategy. It is a performance measure that balances the trade-off between false positives and false negatives to help address the challenge of imbalanced datasets common in anomaly detection.
@@ 4.2.1 Introducing the Efficiency-Accuracy-Complexity Score (EACS)
Model complexity is frequently measured by parameter count, a metric that, while useful,doesn’t fully capture the capabilities of transformer-based models like TransNASTSAD. Despite their high parameter count, these models are highly effective at capturing complex data patterns, a strength that justifies their complexity. Their advanced processing capabilities, facilitated by parallelization, make them adept at handling scenarios that demand high accuracy, even when computational resources are abundant. Therefore, when we report the parameter count of TransNAS-TSAD, it is within the broader context of its sophisticated functionality. Additionally, our evaluation emphasizes not just parameter count but also the F1 score and Efficiency- Accuracy-Complexity Score (EACS), ensuring our assessment considers both the model’s anomaly detection effectiveness and its deployability. While other metrics like ROC/AUC are valuable, our focused approach aligns with the specific objectives of TransNAS-TSAD, aiming for a balanced evaluation of its performance.
[…]
$EACS=w_{a}\times\left(\frac{\text{F1}}{\text{F1}_{\text{max}}}\right)+w_{t}\times\left(1-\frac{\text{T}_{\text{train}}}{\text{T}_{\text{max}}}\right)+w_{p}\times\left(1-\frac{\text{P}_{\text{count}}}{\text{P}_{\text{max}}}\right)$ (11)
where: - $w_{a}$ is the weight for the accuracy component (F1 score), set to 0.4, - $w_{t}$ is the weight for the training time efficiency component, set to 0.4, - $w_{p}$ is the weight for the model complexity (parameter count) component, set to 0.2.
These weights were chosen to reflect the importance of accuracy and training efficiency in practical deployments, where they are often prioritized over the complexity of the model. The F1 score (F1) represents the accuracy of the model, $\text{T}_{\text{train}}$ denotes the training time, and $\text{P}_{\text{count}}$ indicates the parameter count of the model.
@@ 5 Results
Evaluating TransNAS-TSAD: Precision, Recall, and F1 Score Analysis
The evaluation of anomaly detection models, as shown in Table LABEL:Table:3, measures TransNAS-TSAD’s performance across diverse datasets including NAB, UCR, MBA, SMAP, MSL, SWaT, WADI, and SMD, using precision, recall, and F1 scores as benchmarks. TransNAS-TSAD demonstrated high F1 scores in datasets such as NAB, UCR, MBA, and SMAP, benefiting from its advanced data processing techniques that adeptly handle complex patterns in both univariate and multivariate time series. This distinguishes it from models like OmniAnomaly and MSCRED. For the MSL dataset, the TranAD model slightly outperforms TransNAS-TSAD (0.9494 vs 0.9482), likely due to TranAD’s parameters being finely tuned to MSL’s unique characteristics. Despite conducting over 100 trials per dataset, this suggests that even more exhaustive optimization could potentially unlock further improvements for TransNAS-TSAD. Its robust performance in the SWaT dataset, with an F1 score of 0.8314, underscores its versatility in various industrial contexts. In the cases of WADI and SMD, where TransNAS-TSAD achieves F1 scores of 0.8400 and 0.9986 respectively, the significant improvements—such as a 40% increase over baseline in WADI—are attributable to its comprehensive search space and optimization strategies, enabling effective model tuning. This evaluation underscores the potential benefits of extending our optimization framework to achieve even greater model refinement.
The variable performance of models like LSTM-NDT and DAGMM across different datasets points to their methodological strengths and constraints. For example, the LSTM-NDT’s reduced efficacy likely stems from the method’s nonparametric thresholding and resulting sensitivity to large differences in anomaly patterns; and the DAGMM struggles with longer sequences, in part, due to its reliance on a singular GRU model, which limits its capacity to accurately map extended temporal dynamics. Moreover, TransNAS-TSAD demonstrates enhanced performance over innovative yet less adaptable models such as MAD-GAN and USAD, and while CAE-M and GDN show potential in specific scenarios, they generally fall short of matching TransNAS-TSAD’s applicability and performance across the datasets reviewed, highlighting the value of integrating adversarial training and iterative optimization into the TransNAS-TSAD’s framework.
, derived via NSGA-II-based NAS, across diverse datasets.,label=tab:Benchmarking,nofloat=table Method NAB UCR MBA SMAP P R F1 P R F1 P R F1 P R F1 LSTM-NDT 0.6400 0.6667 0.6531 0.5231 0.8294 0.5231 0.9207 0.9718 0.9456 0.8523 0.7326 0.7879 DAGMM 0.7622 0.7292 0.7453 0.5337 0.9718 0.5337 0.9475 0.9900 0.9683 0.8069 0.9891 0.8888 OmniAnomaly 0.8421 0.6667 0.7442 0.8346 0.9999 0.8346 0.8561 1.000 0.9225 0.813 0.9419 0.8728 MSCRED 0.8522 0.6700 0.7502 0.5441 0.9718 0.5441 0.9272 1.0000 0.9623 0.8175 0.9216 0.8664 MAD-GAN 0.8666 0.7012 0.7752 0.8538 0.9891 0.8538 0.9396 1.0000 0.9689 0.8157 0.9216 0.8654 USAD 0.8421 0.6667 0.7442 0.8952 1.0000 0.8952 0.8953 0.9989 0.9443 0.748 0.9627 0.8419 CAE-M 0.7918 0.8019 0.7968 0.6981 1.0000 0.6981 0.8442 0.9997 0.9154 0.8193 0.9567 0.8827 GDN 0.8129 0.7872 0.7998 0.6894 0.9988 0.6894 0.8832 0.9892 0.9332 0.748 0.9891 0.8518 TranAD 0.8889 0.9892 0.9364 0.9407 1.0000 0.9407 0.9569 1.0000 0.978 0.8043 0.9999 0.8915 TransNAS-TSAD 0.8888 1.0000 0.9411 0.9823 1.0000 0.9910 0.9726 1.0000 0.9861 0.9066 1.0000 0.9510 Method MSL SWaT WADI SMD P R F1 P R F1 P R F1 P R F1 LSTM-NDT 0.6288 1.0000 0.7721 0.7778 0.5109 0.6167 0.0138 0.7823 0.0271 0.9736 0.844 0.9042 DAGMM 0.7363 1.0000 0.8482 0.9933 0.6879 0.8128 0.076 0.9981 0.1412 0.9103 0.9914 0.9491 OmniAnomaly 0.7848 0.9924 0.8765 0.9782 0.6957 0.8131 0.3158 0.6541 0.426 0.8881 0.9985 0.9401 MSCRED 0.8912 0.9862 0.9363 0.9992 0.677 0.8072 0.2513 0.7319 0.3741 0.7276 0.9974 0.8414 MAD-GAN 0.8516 0.993 0.9169 0.9593 0.6957 0.8065 0.2233 0.9124 0.3588 0.9991 0.844 0.915 USAD 0.7949 0.9912 0.8822 0.9977 0.6879 0.8143 0.1873 0.8296 0.3056 0.906 0.9974 0.9495 CAE-M 0.7751 1.0000 0.8733 0.9697 0.6957 0.8101 0.2782 0.7918 0.4117 0.9082 0.9671 0.9367 GDN 0.9308 0.9892 0.9591 0.9697 0.6957 0.8101 0.2912 0.7931 0.426 0.717 0.9974 0.8342 TranAD 0.9038 0.9999 0.9494 0.976 0.6997 0.8151 0.3529 0.8296 0.4951 0.9262 0.9974 0.9605 TransNAS-TSAD 0.9567 1.000 0.94.82 0.9415 0.7624 0.8314 0.8508 0.8295 0.8400 0.9985 0.9988 0.9986
@@
In-Depth Analysis of Training Efficiency and EACS in Anomaly Detection Models
Table LABEL:Tab:_EACS presents a comprehensive analysis that compares the Efficiency-Accuracy-Complexity Score (EACS) among various anomaly detection models, including TransNAS-TSAD. It highlights how training time and model size significantly influence the application of anomaly detection models in operational environments that demand rapid deployment and updates, a factor underscored by several studies [29, 55, 17]. In this analysis, TransNAS-TSAD emerges as a standout performer, achieving high EACS values across multiple datasets, which indicates its effective balance of model performance and operational efficiency. Specifically, TransNAS-TSAD delivers high F1 scores of 94.11% and 99.10% on the NAB and UCR datasets, respectively, while maintaining brief training periods of 2.70 and 2.24 seconds. This translates into EACS values of 0.9742 and 0.9922, showcasing TransNAS-TSAD’s rapid adaptability and ability to learn from complex data. The trend (i.e., high accuracies combined with short training times) is consistent across other datasets, including MBA and SMAP, where TransNAS-TSAD records EACS scores of 0.9932 and 0.9734, respectively. This performance sets a stark contrast to other models, such as MSCRED, which achieves commendable F1 scores, but experiences a notable decrease in EACS due to prolonged training times (e.g., 774.99 seconds for MBA). In this light, TransNAS-TSAD distinguishes itself by adeptly optimizing both model complexity and training efficiency, showcasing its ability to balance high-performance metrics with operational practicality. The latter is particularly evident in datasets with challenging environments like SWaT and WADI, where TransNAS-TSAD not only sustains high F1 scores but also significantly reduces training times, achieving EACS scores of 0.9297 and 0.7883, respectively. While models like GDN and TranAD perform well in specific datasets, they do not consistently match TransNAS-TSAD in training efficiency across the board. This evaluation emphasizes the importance of considering training time and model size when deploying practical anomaly detection models.